Explain the significance of Noise Margin? Calculate the Noise Margin for the hardware system following static discipline with voltage levels provided below:
VIL = 1.7 V VIH = 3.2 V VOL = 0.6 V VOH = 4.4 V
1
Expert's answer
2020-11-09T14:18:04-0500

Noise margin: The maximum voltage amplitude of extraneous signal that can be algebraically added to the noise-free worst-case input level without causing the output voltage to deviate from the allowable logic voltage level.

Voltage value between inverter passing from logic maximum (1) to logic minimum (0):"N_{MH}=V_{OH}-V_{IH}=4.4-3.2=1.2" V.

Voltage value between inverter passing from logic minimum (0) to logic maximum (1):

"N_{ML}=V_{IL}-V_{OL}=1.7-0.6=1.1" V.